# HKAN: Hierarchical Kolmogorov-Arnold Network without Backpropagation

### v1

@@ III-C Hyperparameters
HKANs with more layers, more nodes, more BaFs and with smaller values of the parameters $\sigma$ and $\lambda$ tend to fit the target function more accurately. However, such configurations are more susceptible to overfitting. Therefore, these parameters must be carefully tuned to strike an optimal balance between the model’s bias and variance.
@@ III-D Complexity
In layer $l$, each block performs linear regression on $m^{(l)}$ BaFs with a complexity of $O(Nm^{(l)2}+m^{(l)3})$. For layers $l=1,...,L-1$, each containing $n^{(l-1)}n^{(l)}$ blocks, the total complexity per layer is $O(n^{(l-1)}n^{(l)}(Nm^{(l)2}+m^{(l)3}))$. The top layer contains $n^{(L)}$ blocks, resulting in a complexity of $O(n^{(L)}(Nm^{(L)2}+m^{(L)3}))$.
Each function $h^{(l)}_{q}$ linearly combines $n^{(l-1)}$ blocks, yielding a complexity of $O(Nn^{(l-1)2}+n^{(l-1)3})$. For layers $l=1,...,L-1$, with $n^{(l)}$ such functions per layer, the total complexity becomes $O(n^{(l)}(Nn^{(l-1)2}+n^{(l-1)3}))$. The final layer produces a single output, resulting in a complexity of $O(Nn^{(L)2}+n^{(L)3})$.
@@ IV-A Basis Functions
KAN employs B-splines of order 3, which are bell-shaped and computed recursively using the Cox-de Boor formula. The properties of these BaFs, including their number, location, width, and support, are determined by knots. In KANs, these knots are positioned at equidistant intervals, resulting in an even distribution of the basis functions across the input space. The number of knots, which directly influences the spline’s flexibility and the model’s capacity to capture complex patterns, is a crucial hyperparameter in the KAN architecture.
@@ IV-D Learning
Table II summarizes the key differences between KAN and HKAN, providing a quick reference for comparison.
TABLE II: Comparison of KAN and HKAN.
@@
      KAN: Weighted sum of spline (linear combination of BaFs) and SiLU
@@
      KAN: Iterative using backpropagation (gradient descent)
      HKAN: Non-iterative and hierarchical using least-squares method
@@
      HKAN: Importance of inputs can be evaluated based on BlFs
@@ VI Experimental Study
In this section, we compare our proposed HKAN with the standard KAN in regression tasks, evaluating their approximation accuracy. The experimental evaluation was performed on a variety of datasets to validate the model’s performance across different regression and function approximation scenarios.
@@ VI-A Datasets
$g(\mathbf{x})=1-\cos\left(2\pi\sqrt{\sum_{i=1}^{n}x_{i}^{2}}\right)-0.1\sqrt{\sum_{i=1}^{n}x_{i}^{2}},\,n=10,x_{i}\in[-4,4]$
@@
$g(\mathbf{x})=-{\sum_{i=1}^{n}\sin(x_{i})\sin^{20}\left(\frac{ix_{i}^{2}}{\pi}\right)},\,n=2,5,x_{i}\in[0,\pi]$
@@
Figure 3: Synthetic TFs.
All functions, except TF4, have two input arguments. TF4 has ten arguments, while TF5 is evaluated both as a two-argument function and a five-argument variant (TF5-5). The function values and input arguments of TF3, TF4, and TF5 were normalized to the range $[0,1]$. Additionally, the TF2 training data was perturbed by adding noise generated from $U(-0.2,0.2)$.
Table V provides an overview of all datasets used in this study, including six synthetically generated datasets and 18 obtained from various sources, as detailed in the table. For these 18 datasets, both input and output variables were normalized to the range $[0,1]$. The table also specifies the number of samples, input dimensions, and the sizes of the training and test sets. All datasets are available in our GitHub repository [51].
@@ VI-B Optimization
Table III outlines the search space for HKAN hyperparameters. The tree-structured Parzen estimator algorithm, implemented in the Optuna framework [53], was used to explore this space. A total of 1000 trials were conducted, with early stopping applied by pruning trials where the RMSE exceeded twice the baseline RMSE, calculated as the RMSE of the mean prediction on the training set. The optimal hyperparameters were determined using 5-fold cross-validation. The selected hyperparameter values are summarized in Table IV.
TABLE III: Hyperparameter search space for HKAN.
@@
$\quad{}^{**}$nly for the output layer.
TABLE IV: Hyperparameters selected for HKAN.
@@ VI-C Results
Table V summarizes the performance metrics for KAN and HKAN, including the median and interquartile range (IQR) of RMSE for both training and test data, calculated from 50 independent training sessions per model. These results are further illustrated in Fig. 4 using boxplots.
TABLE V: Performance comparison of KAN and HKAN.
@@
Figure 4: Distribution of training (tr) and test (ts) RMSE for KAN and HKAN.
The test errors of both models were compared using the Wilcoxon signed-rank test. As shown in Table V, HKAN achieved significantly lower errors than KAN in 12 out of 24 cases, while KAN outperformed HKAN in 9 cases. Notably, HKAN demonstrated superior accuracy on synthetic functions TF1, TF3, TF5, and TF5-5, where its accuracy exceeded that of KAN by several orders of magnitude. For TF2 (a synthetic function with noise), HKAN’s errors were over 19% lower than KAN’s. A substantial improvement was also observed for the MachineCPU dataset, where HKAN outperformed KAN with a difference exceeding 51%.
@@
When comparing training and test errors for each model separately, it was observed that, in most cases, the training error was significantly lower than the test error, as confirmed by the Wilcoxon test. Cases where the test error was significantly lower than the training error were rare, occurring five times for HKAN and only once for KAN (see the underlined errors in Table V). Substantially lower training errors compared to test errors may indicate overfitting, differences in the distributions of training and test sets, insufficient representation of the test set in the training data, or an inadequate number of training samples relative to the number of input arguments.
The IQR serves as a measure of model variance, reflecting the consistency of predictions across different training sessions. As shown in Table V and Fig. 4, HKAN produces more stable results than KAN in all cases except for the Elevators and Stock datasets. HKAN frequently exhibits an IQR of 0, indicating deterministic behavior. This occurs when the optimal architecture comprises a single layer ($L=1$) with identity BaFs or uniformly spaced BaFs (refer to Table IV for the optimal hyperparameters). Under such conditions, HKAN consistently produces identical results across all training sessions.
@@ VI-D How HKAN Constructs Fitted Function
    Shapes and Complexity of the Fitted Functions: The fitted function evolves in shape and complexity as it progresses through the layers. At the first level, the target function is modeled nonlinearly using individual input variables, capturing only the features apparent in these variables, such as localized fluctuations. The second level integrates these preliminary approximations across all input variables, producing a multi-variable approximation that remains relatively coarse. The subsequent two levels — nonlinear transformations by the blocks of the second layer followed by linear combinations of their outputs — significantly improve the approximation quality. The fifth and sixth levels refine the result by processing the fourth-level outputs in a similar manner.
@@ VI-F Discussion
HKAN eliminates the iterative backpropagation process, significantly reducing computational complexity. By transforming the optimization problem into multiple convex subproblems solved using least-squares regression, HKAN ensures efficient training while maintaining accuracy. Its deterministic training process enhances stability, while the layer-by-layer hierarchical approach facilitates transparent function representation. The flexibility of HKAN’s architecture offers a notable advantage in capturing diverse functional relationships within the data.
@@
HKAN incorporates two ensemble-like principles that enhance its generalization and robustness:
  1. 1.
    Horizontal Integration: Within each layer, diverse blocks are combined, akin to ensemble methods. Each block offers a unique perspective on the input data, and their aggregation enables the network to capture a broader range of patterns and relationships.
  2. 2.
    Vertical Integration: The layer-by-layer processing corresponds to multi-level stacking, where each successive layer builds on the approximations learned by the previous one. This approach allows the network to construct increasingly complex functions.
These ensemble-like properties enable HKAN to model intricate relationships in the data without relying on backpropagation, offering a distinctive approach to function approximation and pattern recognition tasks.
@@
The complexity of the target function directly influences the optimal architecture of HKAN. More complex target functions require deeper and wider networks, while simpler functions can be effectively modeled with fewer layers, blocks, and BaFs. HKAN’s flexibility in selecting BaFs further sets it apart. While KAN relies on B-splines, which can be computationally intensive due to recursive processing, HKAN supports alternative BaFs such as Gaussian and sigmoid, among others. Furthermore, HKAN allows for data-driven or random distributions of BaFs, whereas KAN typically employs evenly distributed BaFs. The smoothing parameter in HKAN is another critical hyperparameter, providing additional control over the model’s adaptability.

### v2

@@ III-C Hyperparameters
HKANs with more layers, more nodes, more BaFs and with smaller values of the parameters $\sigma$ and $\lambda$ tend to fit the target function more accurately. However, such configurations are more susceptible to overfitting. Therefore, these parameters must be tuned to strike an optimal balance between the model’s bias and variance.
@@ III-D Complexity
In layer $l$, each block performs linear regression on $m^{(l)}$ BaFs with a complexity of $O(Nm^{(l)2}+m^{(l)3})$. For layers l=1,,L1l=1,...,L-1, each containing $n^{(l-1)}n^{(l)}$ blocks, the total complexity per layer is $O(n^{(l-1)}n^{(l)}(Nm^{(l)2}+m^{(l)3}))$. The top layer contains $n^{(L)}$ blocks, resulting in a complexity of $O(n^{(L)}(Nm^{(L)2}+m^{(L)3}))$.
Each function $h^{(l)}_{q}$ linearly combines $n^{(l-1)}$ blocks, yielding a complexity of $O(Nn^{(l-1)2}+n^{(l-1)3})$. For layers l=1,,L1l=1,...,L-1, with $n^{(l)}$ such functions per layer, the total complexity becomes $O(n^{(l)}(Nn^{(l-1)2}+n^{(l-1)3}))$. The final layer produces a single output, resulting in a complexity of $O(Nn^{(L)2}+n^{(L)3})$.
@@ IV-A Basis Functions
KAN introduced in [1] employs B-splines of order 3, which are bell-shaped and computed recursively using the Cox-de Boor formula. The properties of these BaFs, including their number, location, width, and support, are determined by knots. In KANs, these knots are positioned at equidistant intervals, resulting in an even distribution of the basis functions across the input space. The number of knots, which directly influences the spline’s flexibility and the model’s capacity to capture complex patterns, is a crucial hyperparameter in the KAN architecture.
@@ IV-C Explainability and Function Representation
Other aspects of HKAN explainability are discussed in Section VI-F. Among these, a particularly valuable feature is the ability to evaluate and interpret the quality of approximation at each stage of hierarchical processing (see Figs. 58).
@@ IV-D Learning
Table III summarizes the key differences between KAN and HKAN, providing a quick reference for comparison.
TABLE III: Comparison of KAN [1] and HKAN.
@@
KAN: Weighted sum of splines (linear combination of BaFs) and SiLU
@@
KAN: Iterative, gradient-based (backpropagation)
HKAN: Non-iterative, hierarchical (least-squares method)
@@
HKAN: Input importance can be assessed based on BlFs; Allows
       insight into approximation quality at each hierarchical level
@@ VI Experimental Study
In this section, we compare our proposed HKAN with the standard KAN and MLP in regression tasks, evaluating their approximation accuracy. The experimental evaluation was performed on a variety of datasets to validate the model’s performance across different regression and function approximation scenarios.
The experiments were performed on a virtual machine equipped with 32 vCPUs (host: 2×AMD EPYC 7302), 64 GB of RAM, running Debian 12. The software stack included Python 3.12.6, PyTorch 2.7.1, NumPy 2.1.1, and scikit-learn 1.5.2.
@@ VI-A Datasets
$g(\mathbf{x})=1-\cos\left(2\pi\sqrt{\sum_{i=1}^{n}x_{i}^{2}}\right)-0.1\sqrt{\sum_{i=1}^{n}x_{i}^{2}},\\ \,n=10,x_{i}\in[-4,4]$
@@
$g(\mathbf{x})=-{\sum_{i=1}^{n}\sin(x_{i})\sin^{20}\left(\frac{ix_{i}^{2}}{\pi}\right)},\,n=2,5,\\ x_{i}\in[0,\pi]$
@@
Figure 3: Synthetic TFs (2D variants of TF4 and TF5).
All functions, except TF4, have two input arguments. TF4 has ten arguments, while TF5 is evaluated both as a two-argument function and a five-argument variant (TF5-5). The function values and input arguments of TF3, TF4, and TF5 were normalized to the range $[0,1]$. Additionally, the TF2 training data was perturbed by adding noise generated from U(0.2,0.2)U(-0.2,0.2).
Table VI provides an overview of all datasets used in this study, including six synthetically generated datasets and 18 obtained from various sources, as detailed in the table. For these 18 datasets, both input and output variables were normalized to the range $[0,1]$. The table also specifies the number of samples, input dimensions, and the sizes of the training and test sets. All datasets are available in our GitHub repository [63].
@@ VI-B Optimization
Table IV outlines the search space for HKAN hyperparameters. The tree-structured Parzen estimator algorithm, implemented in the Optuna framework [65], was used to explore this space. A total of 1000 trials were conducted, with early stopping applied by pruning trials where the RMSE exceeded twice the baseline RMSE, calculated as the RMSE of the mean prediction on the training set. The optimal hyperparameters were determined using 5-fold cross-validation. The selected hyperparameter values are summarized in Table V.
TABLE IV: Hyperparameter search space for HKAN.
@@
[…]
TABLE V: Hyperparameters selected for HKAN.
@@
MLP optimization was carried out using Optuna with 2000 trials. The optimized hyperparameters included depth, hidden size, activation function type, learning rate, batch size, dropout rate, and patience. The search ranges for these hyperparameters were adjusted according to the size of the training datasets. For example, for smaller datasets the maximum depth was limited to 2 and the hidden size to 64, whereas for the largest datasets the depth was increased to 5 and the hidden size to 256.
@@ VI-C Results
Table VI summarizes the performance metrics for KAN and HKAN, including the median and interquartile range (IQR) of RMSE for both training and test data, calculated from 50 independent training sessions per model. These results are further illustrated in Fig. 4 using boxplots.
TABLE VI: Performance comparison of KAN and HKAN.
@@
Figure 4: Distribution of training (tr) and test (ts) RMSE for KAN, MLP and HKAN.
The test errors of both models were compared using the Wilcoxon signed-rank test. As shown in Table VI, HKAN achieved significantly lower errors than KAN in 12 out of 24 cases, while KAN outperformed HKAN in 9 cases. Notably, HKAN demonstrated superior accuracy on synthetic functions TF1, TF3, TF5, and TF5-5, where its accuracy exceeded that of KAN by several orders of magnitude. For TF2 (a synthetic function with noise), HKAN’s errors were over 19% lower than KAN’s. A substantial improvement was also observed for the MachineCPU dataset, where HKAN outperformed KAN with a difference exceeding 51%.
@@
When comparing training and test errors for each model separately, it was observed that, in most cases, the training error was significantly lower than the test error, as confirmed by the Wilcoxon test. Cases where the test error was significantly lower than the training error were rare, occurring five times for HKAN and only once for KAN (see the underlined errors in Table VI). Substantially lower training errors compared to test errors may indicate overfitting, differences in the distributions of training and test sets, insufficient representation of the test set in the training data, or an inadequate number of training samples relative to the number of input arguments.
The IQR serves as a measure of model variance, reflecting the consistency of predictions across different training sessions. As shown in Table VI and Fig. 4, HKAN produces more stable results than KAN in all cases except for the Elevators and Stock datasets. HKAN frequently exhibits an IQR of 0, indicating deterministic behavior. This occurs when the optimal architecture comprises a single layer ($L=1$) with identity BaFs or uniformly spaced BaFs (refer to Table V for the optimal hyperparameters). Under such conditions, HKAN consistently produces identical results across all training sessions.
Table VII compares HKAN with MLP. In this table, the cases where one model’s test error is significantly lower than the competitor’s are highlighted in bold. There are 15 such cases in favor of HKAN and 8 in favor of MLP. Of particular note are the substantially lower errors achieved by HKAN on synthetic datasets, with the exception of TF4. For the remaining datasets, the differences in test errors did not exceed 10%, except for MachineCPU (approximately 22% in favor of HKAN), Kinematics8nm (approximately 105% in favor of MLP), and Elevators (approximately 12% in favor of MLP).
It is also worth noting that the variance of errors in the case of HKAN is significantly smaller than for MLP. Only in four cases does MLP exhibit lower variance than HKAN (see Fig. 4 and the IQR values in Table VII). This indicates that HKAN provides more consistent and stable predictions across independent training sessions.
TABLE VII: Performance comparison of MLP and HKAN.
MLP HKAN
Data Training RMSE Test RMSE Training RMSE Test RMSE
Median IQR Median IQR Median IQR Median IQR
TF1 8.49E-03 4.80E-03 8.60E-03 4.56E-03 2.68E-14 9.62E-15 3.94E-14 2.00E-14
[…]
TF3 3.66E-02 3.50E-03 3.90E-02 3.59E-03 6.37E-06 $8 \ul5.08E$-06 3.85E-07
[…]
TF5 3.33E-02 8.74E-03 3.36E-02 8.60E-03 2.98E-15 4.58E-16 4.56E-15 2.39E-15
TF5-5 7.43E-02 3.78E-03 7.87E-02 3.33E-03 3.40E-09 8.70E-10 3.93E-09 9.88E-10
Abalone 7.39E-02 7.71E-04 7.59E-02 8.08E-04 7.80E-02 1.39E-17 8.00E-02 0
AutoMPG 6.25E-02 4.54E-03 8.19E-02 7.53E-03 6.80E-02 0 8.38E-02 0
[…]
Compactive 2.27E-02 4.23E-04 2.50E-02 9.26E-04 2.24E-02 1.19E-04 2.37E-02 3.94E-04
Concrete 5.87E-02 7.25E-03 6.75E-02 5.38E-03 4.67E-02 1.26E-03 6.12E-02 2.30E-03
Dee 7.63E-02 3.24E-03 1.05E-01 4.65E-03 8.63E-02 1.10E-04 1.03E-01 1.22E-04
Ele2 9.82E-03 8.15E-04 1.07E-02 8.93E-04 7.88E-03 4.78E-07 1.01E-02 9.85E-07
Elevators 2.85E-02 6.85E-04 2.99E-02 6.18E-04 2.85E-02 1.73E-04 3.35E-02 6.34E-04
Kinematics8nm 4.17E-02 1.62E-03 4.93E-02 1.22E-03 9.55E-02 1.15E-03 1.01E-01 1.09E-03
Kinematics32nh 1.12E-01 5.98E-04 1.25E-01 7.40E-04 1.33E-01 0 1.35E-01 0
Laser 1.91E-02 1.14E-03 2.40E-02 2.36E-03 1.46E-02 7.41E-04 2.56E-02 3.62E-03
MachineCPU 4.00E-02 5.56E-03 5.26E-02 1.57E-02 4.13E-02 0 4.13E-02 0
Pumadyn32nh 4.46E-02 3.32E-03 4.72E-02 3.05E-03 4.19E-02 2.08E-04 4.61E-02 6.03E-04
[…]
Stock 2.53E-02 2.67E-03 3.09E-02 2.58E-03 1.63E-02 4.77E-04 2.81E-02 1.75E-03
Treasury 1.12E-02 6.41E-04 1.21E-02 9.85E-04 9.49E-03 6.41E-05 1.11E-02 1.56E-04
Triazines 1.69E-01 8.73E-03 1.86E-01 7.20E-03 1.83E-01 1.14E-02 1.98E-01 7.51E-03
[…]
Table VIII presents the average training times for the optimal configurations of KAN, MLP, and HKAN. It is worth noting that HKAN achieved the shortest training time for 17 datasets, KAN for 4, and MLP for 3.
TABLE VIII: Average training time in seconds.
Data KAN MLP HKAN
TF1 113.09 36.91 8.80
TF2 74.37 8.88 6.44
TF3 333.50 48.18 38.75
TF4 0.56 3.88 0.02
TF5 553.88 27.12 23.17
TF5-5 633.43 93.27 100.62
Abalone 3.86 4.46 0.01
AutoMPG 0.62 0.39 0.01
Bank32nh 2.38 14.33 0.03
Compactive 28.17 86.18 68.69
Concrete 2.20 1.27 3.95
Dee 0.44 0.47 0.02
Ele2 38.43 4.20 3.25
Elevators 69.18 90.33 24.37
Kinematics8nm 82.42 67.03 58.04
Kinematics32nh 12.94 34.67 0.49
Laser 11.95 2.82 0.52
MachineCPU 0.96 0.16 0.01
Pumadyn32nh 71.70 64.01 23.17
Pyramidines 0.17 0.11 12.23
Stock 4.48 5.57 13.34
Treasury 8.25 3.64 3.50
Triazines 4.48 0.13 13.34
Wizmir 8.25 1.32 3.50
@@ VI-D How HKAN Constructs Fitted Function
    Shapes and Complexity of the Fitted Functions: The fitted function evolves in shape and complexity as it progresses through the layers. At the first level, the target function is modeled nonlinearly using individual input variables, capturing only the features apparent in these variables, such as localized fluctuations. The second level integrates these preliminary approximations across all input variables, producing a multi-variable approximation that remains relatively coarse. The subsequent two levels – nonlinear transformations by the blocks of the second layer followed by linear combinations of their outputs – significantly improve the approximation quality. The fifth and sixth levels refine the result by processing the fourth-level outputs in a similar manner.
@@ VI-F Discussion
VI-F1 Learning Process
HKAN eliminates the iterative backpropagation process, in many cases significantly reducing training time compared to KAN and MLP (see Table VIII). By transforming the optimization problem into multiple convex subproblems solved using least-squares regression, HKAN ensures efficient training while maintaining accuracy. Its deterministic training process enhances stability, while the layer-by-layer hierarchical approach facilitates transparent function representation. The flexibility of HKAN’s architecture offers a notable advantage in capturing diverse functional relationships within the data.
VI-F2 Hierarchical Approximation of the Target Function
@@
VI-F3 Built-in Ensembling
HKAN incorporates two ensemble-inspired principles that enhance its generalization ability and robustness:
  • Horizontal integration: Within each layer, diverse blocks are combined, similar to ensemble methods. Each block provides a unique perspective on the input data, and their aggregation enables the network to capture a broader spectrum of patterns and dependencies.
  • Vertical integration: The layer-by-layer processing corresponds to multi-level stacking, where each successive layer builds upon the approximations learned by the previous one. This hierarchical composition allows the network to construct increasingly complex functional representations.
Unlike KAN, which typically employs evenly distributed BaFs [1], HKAN uses data-driven or randomized BaF distributions. This design promotes diversity among blocks and prevents them from converging to identical parameterizations. The blocks can thus be regarded as weak learners that are aggregated through stacking. In this context, inducing diversity among learners is deliberate and crucial for reinforcing the ensemble effect.
These ensemble-like mechanisms enable HKAN to capture intricate dependencies within the data without relying on backpropagation, offering a distinctive and computationally efficient framework for function approximation and pattern recognition tasks.
VI-F4 Flexibility
The complexity of the target function directly determines the optimal architecture of HKAN. More complex target functions require deeper and wider networks, whereas simpler functions can be effectively modeled with fewer layers, blocks, and BaFs. A distinctive feature of HKAN is its flexibility in selecting BaFs. While KAN relies on B-splines [1], which can be computationally demanding due to their recursive nature, HKAN supports alternative basis functions such as Gaussian and sigmoid, among others. The smoothing parameter in HKAN serves as another critical hyperparameter, offering additional control over the model’s flexibility and adaptability.
VI-F5 Built-in Mechanism for Estimating Input Importance
@@
Compared with KAN, HKAN provides a more limited, feature-level interpretability, focusing on the analysis of univariate component functions and input-variable importance rather than symbolic interpretability.
VI-F6 Explainability
HKAN exhibits a high degree of explainability due to its transparent, hierarchical, and modular structure. Unlike end-to-end gradient-based NNs, HKAN decomposes the learning process into a sequence of convex subproblems, each corresponding to an explicit functional transformation. This structure allows each model component to be individually analyzed and interpreted.
Key factors contributing to HKAN’s explainability include:
  • Transparent hierarchical structure: HKAN constructs the target function layer by layer, allowing detailed examination of each stage of approximation (see Figs. 5-8).
  • Univariate blocks and built-in input importance estimation: Eeach block models the relationship between a single input and the output, making input effects easy to interpret. HKAN inherently quantifies the contribution of each input variable using metrics such as $R^{2}$.
  • Deterministic learning process: The absence of stochastic gradient updates ensures reproducible and stable results, enhancing interpretability (assuming fixed BaF parameters).
  • Explicit functional form: HKAN’s output is represented as nested linear combinations of known basis functions, allowing direct tracing of how each component contributes to the final prediction.
VI-F7 Relation to Adaptive and Ensemble-Based Models
HKAN can also be situated within the broader family of adaptive, aggregating, and expanding models. Classical approaches such as Boosting, Additive Trees, and Multivariate Adaptive Regression Splines (MARS) [70] also avoid strict parametric assumptions and rely on flexible function approximation. What these models share with HKAN is the principle of constructing predictive functions from relatively simple components – be they weak learners in Boosting, piecewise trees in Additive Models, or spline basis functions in MARS – and combining them in a structured way to achieve high predictive accuracy.
Despite this high-level similarity, HKAN is fundamentally different from these methods both in architecture and in training procedure. Boosting and Additive Trees rely on sequentially improving ensembles of decision trees using stage-wise optimization, while MARS constructs adaptive spline bases through forward-backward selection. In contrast, HKAN employs randomized basis functions and performs training exclusively through closed-form least-squares regression without iterative optimization. This results in a deterministic, convex learning procedure that clearly distinguishes HKAN from the aforementioned methods.
By combining principles of randomized learning, hierarchical stacking, and functional approximation rooted in the Kolmogorov-Arnold framework, HKAN constitutes a novel contribution that goes beyond existing approaches.